1. Calculating working capital Roost and Sing Corp.s CFO has decided to take a closer look at the firms short-term assets aThe value of Roost and Sing Corp.s working capital is , while its net working capital is The value of Roost and Sing Corp.s

Drop-down options: (number value), (number value), (number value), (1.37x, 0.40x, 2.50x, 1.33x), (increase, decrease, stay same).
Answer #1

1.

Working Capital = Current Assets

Thus,

Working capital = $480,000

2

Net working capital = Current assets - Current liabilities

Thus,

Net working capital = 480000 - 360000 = $120,000

3

Net Operating Working Capital = (Current Assets - Non-operating Current Assets) - (Current liabilities - Non-operating Curren

Non-operating current liabilities = Notes payable = $135,000

Thus,

Net Operating Working Capital = 480.000 - (360,000 - 135,000

Net Operating Working Capital = $255,000

4.

Current ratio = Current assets/current liabilities

thus,

Current ratio = 480000/360000 = 1.33 x

5.

If inventory purchased with long-term debt, it will increase current assets but will not change current liabilities

thus,

Current ratio will INCREASE

Hope this will help, please do comment if you need any further explanation. Your feedback would be highly appreciated.